OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 2x2 - 24x ? (1 points) Answer -6, 0, 4 6, 0, -4 -6, 0, -4 6, 0, 4

OpenStudy (anonymous):

x^3+2x^2-24x = x(x^2+2x-24) this means x = 0; and solving the quadratice equation within brackets viz. x^2+2x-24=0 you get the other two zeroes